# WisPaper: Your AI Scholar Search Engine

### v1

@@ 4.1 Evaluation Setup
We conduct comprehensive evaluations on the two critical components underlying Deep Search (Section 2.1.2): (1) query understanding and criteria generation (Objective 1 in Section 3.1), and (2) paper-criteria matching accuracy (Objective 2 in Section 3.1). These evaluations assess WisModel’s ability to interpret user intent and accurately identify relevant academic literature.
@@ 4.2 Query Understanding and Criteria Generation
Task Description. This experiment evaluates whether WisModel’s generated outputs (search queries and criteria following the format in Appendix Table 7) align with human expert annotations. This corresponds to the first stage of Deep Search’s agent workflow (Section 2.1.2).
@@
Application Recommendations: Based on our findings:
  • Maximum Quality: WisModel achieves unmatched performance, suitable for mission-critical applications.
  • Cost-Efficiency: DeepSeek-V3.2-Exp offers exceptional value as an open-source alternative.
  • Exact Matching: For tasks requiring precise phrase reproduction, WisModel (BLEU 39.8%) significantly outperforms alternatives.
  • Enterprise Deployments: GPT-4o provides balanced performance with comprehensive ecosystem support.
@@ 4.3 Paper-Criteria Matching
Task Description: This experiment evaluates WisModel’s ability to accurately determine whether papers satisfy specified criteria, corresponding to the second stage of Deep Search (Section 2.1.2). Given the 2,777 queries and their 5,879 human-annotated criteria from Section 5.2, we evaluate matching accuracy across the four assessment categories defined in our taxonomy (Appendix Table 10). Domain-expert PhD students labeled each paper-criteria pair using these categories. The complete prompt structure for this task is provided in Appendix Tables 8 and 9.
@@ 5.1 Academic Retrieval and Discovery
Academic search is a cornerstone of scientific research, with mainstream keyword-based systems such as Google Scholar [^7] https://scholar.google.com/ and PubMed [^8] https://pubmed.ncbi.nlm.nih.gov/ supporting global researchers. However, these tools often fall short in handling complex conceptual queries that demand semantic comprehension, as they rely primarily on surface-level text matching rather than deep insight into research intent [8]. To address this limitation, recent studies have explored query reformulation and expansion techniques, yet these approaches still fail to break through the constraints of literal matching [10, 2, 14, 16]. Subsequent advancements have introduced more sophisticated features: Semantic Scholar [^9] https://www.semanticscholar.org/ integrated citation-based metrics and paper recommendation functions, while Connected Papers [^10] https://www.connectedpapers.com/ enabled visual exploration of citation networks. These tools focus on relationship discovery but lack support for comprehensive literature management.
More recently, LLM-powered agents have brought new breakthroughs to academic search. LitSearch [1] and ResearchArena [11] have begun benchmarking LLM capabilities in this domain, verifying their potential to enhance retrieval performance. PaSa developed an autonomous agent that invokes search tools, parses papers, and navigates citation networks to address complex scholarly queries[9]. PaSa outperforms baselines like Google Scholar and ChatGPT in recall metrics. Nevertheless, PaSa remains confined to the search phase and does not address post-retrieval needs such as literature organization and frontier tracking.
@@ 5.2 LLM Agents for Research Assistance
Recent efforts have shifted toward integrating discrete research tasks into cohesive workflows. Specialized systems like ChemCrow [6] and Coscientist [7] automate chemistry experimentation, while virtual laboratory frameworks [19] assist in wet-lab design. However, these tools are domain-specific and do not support cross-disciplinary literature management. End-to-end automation systems such as AI Scientist [13] and ResearchAgent [4] aim to cover ideation, experimentation, and paper writing. While innovative, their fully autonomous design limits user control and positions them as replacements rather than assistants for human researchers. Agent Laboratory [17] expands the automated workflow to include literature review and report writing but prioritizes generating new research outputs over managing existing literature.
@@ 6 Conclusion
In this paper, we introduced WisPaper, an integrated platform that unifies intelligent literature discovery, systematic knowledge management, and personalized frontier tracking. By combining fast keyword retrieval, agent-based deep semantic search, and a structured Library with AI-driven updates, WisPaper provides a closed-loop workflow that supports the full research cycle from exploration to long-term curation. Our large-scale evaluation demonstrates that WisModel achieves state-of-the-art performance in both query interpretation and paper–criteria matching, substantially outperforming strong commercial and open-source baselines. We believe WisPaper offers a practical step toward more efficient, reasoning-centered academic workflows

### v3

@@ 4 Evaluation
We evaluate the two core capabilities underlying Deep Search: (1) Intent Decomposition, (2) Evidence-Backed Validation (§3.1), and additionally (3) end-to-end paper recall on the TaxoBench benchmark [zhang2026deepresearchagentsretrieve], which measures whether the full Deep Search pipeline can retrieve the key papers that domain experts deem essential for a given research topic.
@@ 4.2 Query Understanding and Criteria Generation
Task Description. This experiment evaluates whether WisModel’s generated outputs (search queries and criteria following the format in Appendix Table 7) align with human expert annotations. This corresponds to the first stage of Deep Search’s agent workflow (Section 2.1).
@@ 4.3 Paper-Criteria Matching
Task Description: This experiment evaluates WisModel’s ability to accurately determine whether papers satisfy specified criteria, corresponding to the second stage of Deep Search (Section 2.1). Given the 2,777 queries and their 5,879 human-annotated criteria from Section 5.2, we evaluate matching accuracy across the four assessment categories defined in our taxonomy (Appendix Table 10). Domain-expert PhD students labeled each paper-criteria pair using these categories. The complete prompt structure for this task is provided in Appendix Tables 8 and 9.
@@
4.4 End-to-End Paper Recall
The preceding experiments evaluate Deep Search’s internal components in isolation. This experiment assesses the full pipeline end-to-end: given only a research topic as input, can WisPaper retrieve the core papers that domain experts consider essential?
Benchmark: We adopt TaxoBench [zhang2026deepresearchagentsretrieve], a benchmark designed to evaluate deep research agents on expert-level literature survey tasks. TaxoBench contains 72 high-quality, highly-cited surveys (average citations: 354.5) from 8 computer science subfields, including multimodal learning, reinforcement learning, alignment, and agents. From these surveys, annotators with CS doctoral degrees manually extracted hierarchical taxonomy trees and mapped 3,815 core papers to specific branches. The benchmark’s Deep Research Mode provides only the survey topic as input and measures
Figure 4: End-to-end paper recall (%) on TaxoBench. WisModel achieves 22.26%, surpassing the previous SOTA (O3, 20.92%).
the agent’s ability to retrieve the 3,815 ground-truth papers—a challenging task on which the previous best system (O3) achieves only 20.92% recall.
[…]
Results: Figuer 4 presents the recall results. WisPaper surpasses the previous state of the art. WisPaper achieves 22.26% Recall, exceeding the previous best result of 20.92% set by O3 in TaxoBench’s Deep Research Mode. This is notable because O3 operates as a general-purpose deep research agent with unconstrained web access, whereas WisPaper retrieves from scholar indexed database, demonstrating the effectiveness of structured criteria-based validation.
@@ 5.1 Academic Retrieval and Discovery
Academic search is a cornerstone of scientific research, with mainstream keyword-based systems such as Google Scholar [^7] https://scholar.google.com/ and PubMed [^8] https://pubmed.ncbi.nlm.nih.gov/ supporting global researchers. However, these tools often fall short in handling complex conceptual queries that demand semantic comprehension, as they rely primarily on surface-level text matching rather than deep insight into research intent [https://doi.org/10.1002/jrsm.1378]. To address this limitation, recent studies have explored query reformulation and expansion techniques, yet these approaches still fail to break through the constraints of literal matching [jagerman2023queryexpansionpromptinglarge, 10.1145/3539618.3591960, ma-etal-2023-query, 10.1145/3589335.3648298]. Subsequent advancements have introduced more sophisticated features: Semantic Scholar [^9] https://www.semanticscholar.org/ integrated citation-based metrics and paper recommendation functions, while Connected Papers [^10] https://www.connectedpapers.com/ enabled visual exploration of citation networks. These tools focus on relationship discovery but lack support for comprehensive literature management.
More recently, LLM-powered agents have brought new breakthroughs to academic search. LitSearch [ajith-etal-2024-litsearch] and ResearchArena [kang-xiong-2025-researcharena] have begun benchmarking LLM capabilities in this domain, verifying their potential to enhance retrieval performance. PaSa developed an autonomous agent that invokes search tools, parses papers, and navigates citation networks to address complex scholarly queries[he-etal-2025-pasa]. PaSa outperforms baselines like Google Scholar and ChatGPT in recall metrics. Nevertheless, PaSa remains confined to the search phase and does not address post-retrieval needs such as literature organization and frontier tracking.
@@ 5.2 LLM Agents for Research Assistance
Recent efforts have shifted toward integrating discrete research tasks into cohesive workflows. Specialized systems like ChemCrow [bran2023chemcrowaugmentinglargelanguagemodels] and Coscientist [gottweis2025aicoscientist] automate chemistry experimentation, while virtual laboratory frameworks [Swanson2024.11.11.623004] assist in wet-lab design. However, these tools are domain-specific and do not support cross-disciplinary literature management. End-to-end automation systems such as AI Scientist [lu2024aiscientistfullyautomated] and ResearchAgent [baek-etal-2025-researchagent] aim to cover ideation, experimentation, and paper writing. While innovative, their fully autonomous design limits user control and positions them as replacements rather than assistants for human researchers. Agent Laboratory [schmidgall2025agentlaboratoryusingllm] expands the automated workflow to include literature review and report writing but prioritizes generating new research outputs over managing existing literature.
@@ 6 Conclusion
We presented WisPaper, a scholarly agent system designed to bridge the gap between information retrieval and knowledge management. By closing the loop between search, organization, and monitoring, WisPaper transforms passive literature consumption into an active, self-refining workflow. Central to our system is WisPaper, an agent explicitly trained to perform structured intent decomposition and rigorous evidence-backed verification. Our evaluation highlights the system’s robustness: WisPaper achieves 93.70% accuracy in paper validation, and the full pipeline attains a recall of 22.26% on the TaxoBench benchmark, surpassing existing state-of-the-art agents. WisPaper is publicly deployed at https://wispaper.ai , serving as a transparent and efficient tool for researchers across disciplines.